# Federated Learning: A new frontier in the exploration of multi-institutional medical imaging data

### v1

@@ 4.5 Label deficiency
Training DL models requires reliable and large labeled datasets to achieve satisfactory results. Frequently, the institutions involved cover only partially annotated or even entirely unannotated datasets. This is not surprising, as the labeling of medical data is very time-consuming and requires sufficient knowledge and experience, making it challenging to construct necessarily labeled datasets [Dercksen et al., 2019, Ji et al., 2023]. In an FL scenario, different clients may have data labeled to varying degrees. In this context, the standard FL scheme (described in section 3.1) cannot be used as it requires all clients to hold fully labeled data. To address this problem, semi-supervised [Mushtaq et al., 2023, Liu et al., 2023b, Wang et al., 2024] and self-supervised learning [Park et al., 2024] have been introduced.
@@ 4.6 System architecture
Another critical factor in FL-based systems is the appropriate computing architecture chosen to ensure the feasibility, effectiveness, and efficiency [Darzidehkalani et al., 2022]. The architecture for a specific MI problem should address the issues raised in previous paragraphs while also controlling client synchronization problems [Aouedi et al., 2022]. More information on optimizing client synchronization will be discussed in section 4.7, as the communication efficiency problem relates to the system architecture. Generally, system architecture covers computing power that can be either private or cloud-based train models locally, storage systems (i.e., local storage or in the cloud), and software employed to archive and maintain the data such as the Picture Archiving and Communication System (PACS) and Radiological Information System (RIS) [Darzidehkalani et al., 2022, Jacobs et al., 2023]. A complete system architecture for the FL partnership should include additional elements that allow tracking of the model training process, including recording data access history, training structure, selection and modification of configuration, and hyperparameter tuning [Kandati and Gadekallu, 2023a]. Such a tracking system is necessary to meet traceability and accountability requirements, as it assists researchers in interpreting the global model by examining the data source. Finally, the tracking system enables one to assess the contribution level of individual participants [Naz et al., 2022].
@@ 4.7 Communication efficiency
To allow institutions to train the model smoothly, a reliable way of communicating with each other must be secured. Communication involves constantly sharing the parameters of global and local models directly between clients or between clients and the global server. This leads to a bottleneck problem if the ability to disseminate information due to network bandwidth is limited [Patel et al., 2022, Aouedi et al., 2022].
@@ 5 Aggregation methods in medical imaging
In MI the most commonly used aggregation method is the standard FedAvg [Roy et al., 2019, Li et al., 2019, Roth et al., 2020, Cetinkaya et al., 2021, Parekh et al., 2021, Ziller et al., 2021b, Liu et al., 2021a, Dou et al., 2021, Feki et al., 2021, Stripelis et al., 2021a, Roth et al., 2021, Stripelis et al., 2021b, Agbley et al., 2021, Linardos et al., 2022, Adnan et al., 2022, Li et al., 2022c, He et al., 2022, Yang et al., 2022b, Subramanian et al., 2022, Luo and Wu, 2022, Ślazyk et al., 2022, Zhou et al., 2022, Elshabrawy et al., 2022, Pati et al., 2022, Misonne and Jodogne, 2022, Stripelis et al., 2022, Lu et al., 2022b, Kumar et al., 2022, Tan et al., 2023, Liu et al., 2023b, Jiménez-Sánchez et al., 2023, Mushtaq et al., 2023, Elmas et al., 2023, Levac et al., 2023, Denissen et al., 2023, Kanhere et al., 2023, Kaushal et al., 2023, Makkar and Santosh, 2023, Wu et al., 2023a, Wang et al., 2023, Kim et al., 2024, Qi et al., 2024b, AlSalman et al., 2024, Mitrovska et al., 2024, Yamada et al., 2024, Yan et al., 2024, Xiang et al., 2024, Zheng et al., 2024, Zhou et al., 2024, Hossain et al., 2024, Sun et al., 2024, Khan et al., 2024, Deng et al., 2024a, Babar et al., 2024, Myrzashova et al., 2025, Gupta et al., 2024b, Albalawi et al., 2024, Gupta et al., 2024a, Kumar et al., 2024, Deng et al., 2024b] or its modification with equal weights (FL-EV) [Li et al., 2020b, Guo et al., 2021, Lo et al., 2021, Florescu et al., 2022, Linardos et al., 2022, Peng et al., 2023, Naumova et al., 2024, Abbas et al., 2024, Liu et al., 2024a, Vo et al., 2024]. Some of the algorithms also use FedProx [Elshabrawy et al., 2022, Subramanian et al., 2022, Qi et al., 2024b], FedBN [Elshabrawy et al., 2022, Kanhere et al., 2023, Kulkarni et al., 2023] or adaptive algorithms [Stripelis et al., 2022, Levac et al., 2023, Qi et al., 2024b]. Other techniques dedicated to MI are modifications or extensions of FedAvg, involving adjusting the weighting factor [Shen et al., 2021, Mächler et al., 2022, Khan et al., 2022, Alkhunaizi et al., 2022, Mächler et al., 2023, Khan et al., 2023, Wicaksana et al., 2023, Wu et al., 2023b, Liu et al., 2023a, Yang et al., 2023, Noman et al., 2023, Lai et al., 2024, Zhang et al., 2024, Wang et al., 2024] or performing the aggregation only on specific parameters of local models, such as those from particular layers, instead of on all parameters [Gunesli et al., 2021, Bernecker et al., 2022, Liu et al., 2023a].
New aggregation techniques aim to overcome the fundamental limitations of FL. The following subsections present the most important features of different aggregation methods categorized into algorithms that enhance the global model accuracy, handle malicious clients, and improve data security.
@@ 5.1 Improving model’s accuracy
Local models trained on small or low-quality datasets can negatively impact the accuracy of the global model, leading to decreased accuracy across all clients participating in federated learning. By introducing new aggregation algorithms, we expect to increase the accuracy of the global model while preventing the accuracy reduction locally. The widely used approach is to dynamically assign weighting factor to each involved client based on their local loss functions [Shen et al., 2021, Mächler et al., 2022, 2023, Wicaksana et al., 2023, Wu et al., 2023b, Liu et al., 2023a]. The methods presented in section 3 use constant weights across global rounds. This section reviews up-to-date methods that employ round-dependent weights.
@@
with $w_{c}^{r}$ being the weight associated with client $c$–th at round $r$–th. The symbol $T$ in Eq. (10) represents a temperature, $\xi$ is used to control the impact of the weights, $\rho_{c}^{r-1}$ is the dynamic proportion of the loss function $\mathcal{L}$ changes from two consecutive rounds $r-1$ and $r-2$.
@@
The second method that nonlinearly transforms the loss function is the Model-based Federated learning (ModFed) [Wu et al., 2023b], which uses the exponential function, as follows:
@@
$w_{c}^{r}=\frac{y_{c}^{r}\cdot n_{c}}{\sum_{{c^{\prime}}=1}^{C}y_{c^{\prime}}^{r}\cdot n_{c^{\prime}}}.$ (19)
@@ 5.3 Security
$\text{Clip}\left(\mathbf{\Theta},D^{r}\right)=\mathbf{\Theta}/\text{max}\left(1,\frac{||\mathbf{\Theta}||_{2}}{D^{r}}\right).$ (26)
@@
with $g=0,...,n_{\epsilon}-1$ being a consensus round, $\eta_{c}$ a set of neighbours connected to client $c$–th that can exchange local model parameters, and $n_{c}$ the number of samples from collaborator $c$, and $\epsilon$ denotes the step size. At the end of the communication, the proxy variables $\mathbf{z}_{c}^{n_{\epsilon}-1}$ approximate the parameters of the global averaged model $\mathbf{\Theta}^{r}$ (cf. Eq. (3)):
@@ 6 Learning methods in medical imaging
Centralized approach is the most common learning method used in context of MI [Li et al., 2019, Roth et al., 2020, Li et al., 2020b, Sheller et al., 2020, Wang et al., 2020c, Liu et al., 2021a, Lo et al., 2021, Zhang et al., 2021, Ziller et al., 2021b, Huang et al., 2021, Lee et al., 2021, Sarma et al., 2021, Gunesli et al., 2021, Agbley et al., 2021, Ke et al., 2021, Cetinkaya et al., 2021, Shen et al., 2021, Tian et al., 2021, Stripelis et al., 2021a, b, Pati et al., 2021, Parekh et al., 2021, Feki et al., 2021, Roth et al., 2021, Malekzadeh et al., 2021, Guo et al., 2021, Dou et al., 2021, Linardos et al., 2022, Bernecker et al., 2022, Ślazyk et al., 2022, Subramanian et al., 2022, Luo and Wu, 2022, Zhang et al., 2022a, Li et al., 2022c, He et al., 2022, Yang et al., 2022a, Alkhunaizi et al., 2022, Zhou et al., 2022, Islam et al., 2022, Lu and Kalpathy-Cramer, 2022, Adnan et al., 2022, Florescu et al., 2022, Elshabrawy et al., 2022, Yang et al., 2022b, Bercea et al., 2022, Guo et al., 2022, Khan et al., 2022, Mächler et al., 2022, Dalmaz et al., 2022, Misonne and Jodogne, 2022, Muthukrishnan et al., 2022, Kumar et al., 2022, Stripelis et al., 2022, Lu et al., 2022b, Baid et al., 2022, Dalmaz et al., 2024, Kandati and Gadekallu, 2023b, Lin et al., 2023, Mächler et al., 2023, Khan et al., 2023, Wicaksana et al., 2023, Tan et al., 2023, Wu et al., 2023b, Kulkarni et al., 2023, Liu et al., 2023a, b, Jiménez-Sánchez et al., 2023, Jiang et al., 2023, Hosseini et al., 2023, Yang et al., 2023, Zhao and Huang, 2023, Mushtaq et al., 2023, Feng et al., 2023, Zhou et al., 2023, Elmas et al., 2023, Levac et al., 2023, Rajagopal et al., 2023, Denissen et al., 2023, Makkar and Santosh, 2023, Peng et al., 2023, Kaushal et al., 2023, Wu et al., 2023a, Wang et al., 2023, Kim et al., 2024, Naumova et al., 2024, Qi et al., 2024b, AlSalman et al., 2024, Mitrovska et al., 2024, Yamada et al., 2024, Yan et al., 2024, Lai et al., 2024, Xiang et al., 2024, Abbas et al., 2024, Zheng et al., 2024, Liu et al., 2024a, Zhou et al., 2024, Hossain et al., 2024, Sun et al., 2024, Deng et al., 2024a, Babar et al., 2024, Ahmed et al., 2024, Kong et al., 2024, Shiri et al., 2024, Gupta et al., 2024b, Zhang et al., 2024, Albalawi et al., 2024, Vo et al., 2024, Liang et al., 2024, Wang et al., 2024, Deng et al., 2024b], although a decentralized federated learning, peer-to-peer [Roy et al., 2019, Wu et al., 2021, Giuseppi et al., 2022, Huang et al., 2022, Kalra et al., 2023, Khan et al., 2024] or blockchain-based approach [Nguyen et al., 2022a, Kalapaaking et al., 2023, Noman et al., 2023, Myrzashova et al., 2025, Gupta et al., 2024a, Kumar et al., 2024], can also be found in some articles.
@@
In recent years, several modern solutions have emerged to overcome the significant challenges of federated learning, such as data heterogeneity [Li et al., 2020b, Guo et al., 2021, Liu et al., 2021a, Guo et al., 2021, Huang et al., 2021, Yang et al., 2022a, Bercea et al., 2022, Zhu and Luo, 2022, Zhang et al., 2022a, Dalmaz et al., 2022, 2024, Luo and Wu, 2022, Zhao and Huang, 2023, Elmas et al., 2023, Feng et al., 2023, Wu et al., 2023b, Jiang et al., 2023, Jiménez-Sánchez et al., 2023, Yang et al., 2023, Lin et al., 2023], or to improve the data security [Li et al., 2019, Ziller et al., 2021b, Malekzadeh et al., 2021, Yang et al., 2022b, Nguyen et al., 2022a, Kalra et al., 2023]. This section presents the learning algorithms addressing limitations specific to MI, including model accuracy, data heterogeneity, label deficiency, and security.
@@ 6.1 Improving model accuracy
An alternative approach used to improve the accuracy of a global model is to follow continual learning [Sheller et al., 2020] with a particular focus on the use of synaptic intelligence (SI) [Huang et al., 2022]. Synapses accumulate task-relevant information over time, enabling them to efficiently create and retain new memories without forgetting old ones. This technique has been successfully applied for segmenting brain metastases using structural MRI data. An example of SI is the Continual Learning for Peer-to-Peer Federated Learning approach [Huang et al., 2022]. This method introduces a slightly different peer-to-peer scheme than the one discussed in section 3.4. The clients do not share their trained local models for the aggregation procedure. Instead, a single model is trained and transferred successively from one institution to another, repeating this process multiple times. This approach penalizes alternations in model parameters that are important for gaining knowledge from the previous client’s data.
@@
The RoS-FL has already shown robust results in the segmentation and restoration of medical images [Yang et al., 2022b].

### v2

@@ 4.5 Label deficiency
Training DL models requires reliable and large labeled datasets to achieve satisfactory results. Frequently, the institutions involved cover only partially annotated or even entirely unannotated datasets. This is not surprising, as the labeling of medical data is very time-consuming and requires sufficient knowledge and experience, making it challenging to construct necessarily labeled datasets [Dercksen et al., 2019, Ji et al., 2024]. In an FL scenario, different clients may have data labeled to varying degrees. In this context, the standard FL scheme (described in section 3.1) cannot be used as it requires all clients to hold fully labeled data. To address this problem, semi-supervised [Mushtaq et al., 2023, Liu et al., 2023b, Wang et al., 2024] and self-supervised learning [Park et al., 2024] have been introduced.
@@ 4.6 System architecture
Another critical factor in FL-based systems is the appropriate computing architecture chosen to ensure the feasibility, effectiveness, and efficiency [Darzidehkalani et al., 2022b]. The architecture for a specific MI problem should address the issues raised in previous paragraphs while also controlling client synchronization problems [Aouedi et al., 2022]. More information on optimizing client synchronization will be discussed in section 4.8, as the communication efficiency problem relates to the system architecture. Generally, system architecture covers computing power that can be either private or cloud-based train models locally, storage systems (i.e., local storage or in the cloud), and software employed to archive and maintain the data such as the Picture Archiving and Communication System (PACS) and Radiological Information System (RIS) [Darzidehkalani et al., 2022b, Jacobs et al., 2023]. A complete system architecture for the FL partnership should include additional elements that allow tracking of the model training process, including recording data access history, training structure, selection and modification of configuration, and hyperparameter tuning [Kandati and Gadekallu, 2023a]. Such a tracking system is necessary to meet traceability and accountability requirements, as it assists researchers in interpreting the global model by examining the data source. Finally, the tracking system enables one to assess the contribution level of individual participants [Naz et al., 2022].
4.7 Model heterogeneity
One of the key assumptions of the standard FL is that all clients participating in the training process use the same neural network architecture [Alam et al., 2022]. However, this assumption often does not hold due to variations in system architecture, particularly in computational resources. Additionally, different institutions may have specific prerequisites that require diverse model architectures [Nezhad et al., 2025, Weng et al., 2025]. For instance, a client may prefer traditional convolutional networks for their spatial precision and computational efficiency. In contrast, others may choose transformers to effectively capture long-range dependencies and enable context-aware processing [Nezhad et al., 2025]. Some institutions may also use other models, such as random forests [Afonin and Karimireddy, 2022]. In these scenarios, employing standard FL becomes impractical, and requiring all institutions to use the same model architecture could be inefficient and hinder the performance [Nezhad et al., 2025, Weng et al., 2025].
@@ 4.8 Communication efficiency
To allow institutions to train the model smoothly, a reliable way of communicating with each other must be secured. Communication involves constantly sharing the parameters of global and local models directly between clients or between clients and the global server. This leads to a bottleneck problem if the ability to disseminate information due to network bandwidth is limited [Patel et al., 2022, Aouedi et al., 2022]. This issue is especially significant for medical imaging models, which often require high-capacity architectures to process high-resolution volumetric data. Consequently, the communication overhead becomes a practical constraint in hospital networks that have limited or regulated bandwidth.
@@ 5 Aggregation methods in medical imaging
In MI the most commonly used aggregation method is the standard FedAvg [Roy et al., 2019, Li et al., 2019, Roth et al., 2020, Cetinkaya et al., 2021, Parekh et al., 2021, Ziller et al., 2021b, Liu et al., 2021a, Dou et al., 2021, Feki et al., 2021, Stripelis et al., 2021a, Roth et al., 2021, Stripelis et al., 2021b, Linardos et al., 2022, Adnan et al., 2022, Li et al., 2022c, He et al., 2022, Yang et al., 2022b, Subramanian et al., 2022, Luo and Wu, 2022, Ślazyk et al., 2022, Zhou et al., 2022, Elshabrawy et al., 2022, Pati et al., 2022, Misonne and Jodogne, 2022, Stripelis et al., 2022, Lu et al., 2022b, Kumar et al., 2022, Tan et al., 2023, Liu et al., 2023b, Jiménez-Sánchez et al., 2023, Mushtaq et al., 2023, Elmas et al., 2023, Levac et al., 2023, Denissen et al., 2023, Kanhere et al., 2023, Kaushal et al., 2023, Makkar and Santosh, 2023, Wu et al., 2023, Wang et al., 2023, Kim et al., 2024, Qi et al., 2024b, AlSalman et al., 2024, Mitrovska et al., 2024, Yamada et al., 2024, Yan et al., 2024, Xiang et al., 2024, Zheng et al., 2024, Zhou et al., 2024, Hossain et al., 2024, Sun et al., 2024, Khan et al., 2024, Deng et al., 2024a, Babar et al., 2024, Myrzashova et al., 2025, Gupta et al., 2024b, Albalawi et al., 2024, Gupta et al., 2024a, Kumar et al., 2024, Deng et al., 2024b] or its modification with equal weights (FL-EV) [Li et al., 2020b, Guo et al., 2021, Lo et al., 2021, Florescu et al., 2022, Linardos et al., 2022, Peng et al., 2023, Naumova et al., 2024, Abbas et al., 2024, Liu et al., 2024a, Vo et al., 2024]. Some of the algorithms also use FedProx [Elshabrawy et al., 2022, Subramanian et al., 2022, Qi et al., 2024b], FedBN [Elshabrawy et al., 2022, Kanhere et al., 2023, Kulkarni et al., 2023] or adaptive algorithms [Stripelis et al., 2022, Levac et al., 2023, Qi et al., 2024b]. Other techniques dedicated to MI are modifications or extensions of FedAvg, involving adjusting the weighting factor [Shen et al., 2021, Mächler et al., 2022, Khan et al., 2022, Alkhunaizi et al., 2022, Mächler et al., 2023, Khan et al., 2023, Wicaksana et al., 2023, Wu et al., 2024, Liu et al., 2023a, Yang et al., 2023, Noman et al., 2023, Lai et al., 2024, Zhang et al., 2024, Wang et al., 2024] or performing the aggregation only on specific parameters of local models, such as those from particular layers, instead of on all parameters [Gunesli et al., 2021, Bernecker et al., 2022, Liu et al., 2023a].
New aggregation techniques aim to overcome the fundamental limitations of FL. The following subsections present the most important features of different aggregation methods categorized into algorithms that enhance the global model accuracy, handle malicious clients, and improve data security. These strategies operate at the level of model parameters or updates, indicating they are largely independent of specific downstream tasks, such as classification or segmentation, or network architectures. Although individual studies validate these strategies for specific MI tasks, they are generally applicable to a wide range of applications.
@@ 5.1 Improving model’s accuracy
Local models trained on small or low-quality datasets can negatively impact the accuracy of the global model, leading to decreased accuracy across all clients participating in federated learning. By introducing new aggregation algorithms, we expect to increase the accuracy of the global model while preventing the accuracy reduction locally. The widely used approach is to dynamically assign weighting factor to each involved client based on their local loss functions [Shen et al., 2021, Mächler et al., 2022, 2023, Wicaksana et al., 2023, Wu et al., 2024, Liu et al., 2023a]. The methods presented in section 3 use constant weights across global rounds. This section reviews up-to-date methods that employ round-dependent weights.
@@
with $w_{c}^{r}$ being the weight associated with client cc–th at round rr–th. The symbol $T$ in Eq. (10) represents a temperature, $\xi$ is used to control the impact of the weights, $\rho_{c}^{r-1}$ is the dynamic proportion of the loss function $\mathcal{L}$ changes from two consecutive rounds r1r-1 and r2r-2.
@@
The second method that nonlinearly transforms the loss function is the Model-based Federated learning (ModFed) [Wu et al., 2024], which uses the exponential function, as follows:
@@
$w_{c}^{r}=\frac{y_{c}^{r}n_{c}}{\sum_{{c^{\prime}}=1}^{C}y_{c^{\prime}}^{r}n_{c^{\prime}}}.$ (19)
@@ 5.3 Security
where $\text{DP}(.)$ is the function that adds Gaussian noise to the model update, and $\text{Clip}(.)$ is the adaptive quantile clipping function, defined as follows:
@@
with $g=0,...,n_{\epsilon}-1$ being a consensus round, $\eta_{c}$ a set of neighbours connected to client cc–th that can exchange local model parameters, and $n_{c}$ the number of samples from collaborator $c$, and $\epsilon$ denotes the step size. At the end of the communication, the proxy variables $\mathbf{z}_{c}^{n_{\epsilon}-1}$ approximate the parameters of the global averaged model $\mathbf{\Theta}^{r}$ (cf. Eq. (3)):
@@ 6 Learning methods in medical imaging
Centralized approach is the most common learning method used in context of MI [Li et al., 2019, Roth et al., 2020, Li et al., 2020b, Sheller et al., 2020, Wang et al., 2020c, Liu et al., 2021a, Lo et al., 2021, Zhang et al., 2021, Ziller et al., 2021b, Huang et al., 2021, Lee et al., 2021, Sarma et al., 2021, Gunesli et al., 2021, Ke et al., 2021, Cetinkaya et al., 2021, Shen et al., 2021, Tian et al., 2021, Stripelis et al., 2021a, b, Pati et al., 2021, Parekh et al., 2021, Feki et al., 2021, Roth et al., 2021, Malekzadeh et al., 2021, Guo et al., 2021, Dou et al., 2021, Linardos et al., 2022, Bernecker et al., 2022, Ślazyk et al., 2022, Subramanian et al., 2022, Luo and Wu, 2022, Zhang et al., 2022a, Li et al., 2022c, He et al., 2022, Yang et al., 2022a, Alkhunaizi et al., 2022, Zhou et al., 2022, Islam et al., 2022, Lu and Kalpathy-Cramer, 2022, Adnan et al., 2022, Florescu et al., 2022, Elshabrawy et al., 2022, Yang et al., 2022b, Bercea et al., 2022, Guo et al., 2022, Khan et al., 2022, Mächler et al., 2022, Dalmaz et al., 2022, Misonne and Jodogne, 2022, Muthukrishnan et al., 2022, Kumar et al., 2022, Stripelis et al., 2022, Lu et al., 2022b, Baid et al., 2022, Dalmaz et al., 2024, Kandati and Gadekallu, 2023b, Lin et al., 2023, Mächler et al., 2023, Khan et al., 2023, Wicaksana et al., 2023, Tan et al., 2023, Wu et al., 2024, Kulkarni et al., 2023, Liu et al., 2023a, b, Jiménez-Sánchez et al., 2023, Jiang et al., 2023, Hosseini et al., 2023, Yang et al., 2023, Zhao and Huang, 2023, Mushtaq et al., 2023, Feng et al., 2023, Zhou et al., 2023, Elmas et al., 2023, Levac et al., 2023, Rajagopal et al., 2023, Denissen et al., 2023, Makkar and Santosh, 2023, Peng et al., 2023, Kaushal et al., 2023, Wu et al., 2023, Wang et al., 2023, Kim et al., 2024, Naumova et al., 2024, Qi et al., 2024b, AlSalman et al., 2024, Mitrovska et al., 2024, Yamada et al., 2024, Yan et al., 2024, Lai et al., 2024, Xiang et al., 2024, Abbas et al., 2024, Zheng et al., 2024, Liu et al., 2024a, Zhou et al., 2024, Hossain et al., 2024, Sun et al., 2024, Deng et al., 2024a, Babar et al., 2024, Ahmed et al., 2024, Kong et al., 2024, Shiri et al., 2024, Gupta et al., 2024b, Zhang et al., 2024, Albalawi et al., 2024, Vo et al., 2024, Liang et al., 2024, Wang et al., 2024, Deng et al., 2024b], although a decentralized federated learning, peer-to-peer [Roy et al., 2019, Wu et al., 2021, Giuseppi et al., 2022, Huang et al., 2022, Kalra et al., 2023, Khan et al., 2024] or blockchain-based approach [Nguyen et al., 2022a, Kalapaaking et al., 2023, Noman et al., 2023, Myrzashova et al., 2025, Gupta et al., 2024a, Kumar et al., 2024], can also be found in some articles.
@@
In recent years, several modern solutions have emerged to overcome the significant challenges of federated learning, such as data heterogeneity [Li et al., 2020b, Guo et al., 2021, Liu et al., 2021a, Guo et al., 2021, Huang et al., 2021, Yang et al., 2022a, Bercea et al., 2022, Zhu and Luo, 2022, Zhang et al., 2022a, Dalmaz et al., 2022, 2024, Luo and Wu, 2022, Zhao and Huang, 2023, Elmas et al., 2023, Feng et al., 2023, Wu et al., 2024, Jiang et al., 2023, Jiménez-Sánchez et al., 2023, Yang et al., 2023, Lin et al., 2023], or to improve the data security [Li et al., 2019, Ziller et al., 2021b, Malekzadeh et al., 2021, Yang et al., 2022b, Nguyen et al., 2022a, Kalra et al., 2023]. This section presents the learning algorithms addressing limitations specific to MI, including model accuracy, data heterogeneity, label deficiency, and security.
@@ 6.1 Improving model accuracy
An alternative approach used to improve the accuracy of a global model is to follow continual learning [Sheller et al., 2020] with a particular focus on the use of synaptic intelligence (SI) [Huang et al., 2022]. Synapses accumulate task-relevant information over time, enabling them to efficiently create and retain new memories without forgetting old ones. An example of SI is the Continual Learning for Peer-to-Peer Federated Learning approach [Huang et al., 2022]. This method introduces a slightly different peer-to-peer scheme than the one discussed in section 3.4. The clients do not share their trained local models for the aggregation procedure. Instead, a single model is trained and transferred successively from one institution to another, repeating this process multiple times. This approach penalizes alternations in model parameters that are important for gaining knowledge from the previous client’s data.